OpenStudy (ttop0816):

please help!!!Solve sec^2θ + tanθ - 3 = 0 for 0° ≤ θ < 90°. θ = ???

OpenStudy (anonymous):

I think theta = 0 to 90degrees

OpenStudy (anonymous):

plot | -3+sec^2(theta)+tan(theta) = 0 | theta = 0 to 90 °

OpenStudy (ttop0816):

then would it be 45 degrees?? i got that 0.0

OpenStudy (mertsj):

\[1+\tan ^2\theta +\tan \theta -3=0\] \[\tan ^2\theta+\tan \theta -2=0\] \[(\tan \theta +2)(\tan \theta-1)=0\] \[\tan \theta=-2 or \tan \theta=1\] \[\theta=45^{o}\]